# Evaluating The Explainability of State-of-the-Art Machine Learning-based IoT Network Intrusion Detection Systems

### v1

@@ Abstract
Internet-of-Things (IoT) Network Intrusion Detection Systems (NIDSs) which use machine learning (ML) models achieve high detection performance and accuracy while avoiding dependence on fixed signatures extracted from attack artifacts. However, there is a noticeable hesitance among network security experts and practitioners when it comes to deploying ML-based NIDSs in real-world production environments due to their black-box nature, i.e., how and why the underlying models make their decisions. In this work, we analyze state-of-the-art ML-based IoT NIDS models using explainable AI (xAI) techniques (e.g., TRUSTEE, SHAP). Using the explanations generated for the models’ decisions, the most prominent features used by each NIDS model considered are presented. We compare the explanations generated across xAI methods for a given NIDS model as well as the explanations generated across the NIDS models for a given xAI method. Finally, we evaluate the vulnerability of each NIDS model to inductive bias (artifacts learnt from training data). The results show that: (1) some ML-based IoT NIDS models can be better explained than other models, (2) xAI explanations are in conflict for most of the IoT NIDS models considered in this work and (3) some IoT NIDS models are more vulnerable to inductive bias than other models.
@@
Network Intrusion Detection Systems, NIDS, Internet-of-Things, IoT, Machine Learning, explainable Artificial Intelligence, xAI
@@ I Introduction
The Internet-of-Things (IoT) [1] constitutes a vast network of sensor-equipped devices, often characterized by low power and modest processing capabilities. These devices seamlessly exchange data among themselves or with central systems such as gateways or cloud servers using a mix of wired (such as Ethernet) and wireless technologies (such as RFID, Zigbee, WiFi, Bluetooth, and 3G/4G). The applications of IoT devices span a wide spectrum: from wearables like fitness trackers and smartwatches, to home automation systems managing lighting, thermostats, cameras, and door locks, all the way to industrial automation for tasks like manufacturing equipment management, process control, and ensuring plant safety. By 2032, experts estimate that a staggering 34.4 billion IoT devices will be deployed globally [2]. Unfortunately, cyber attackers have set their sights on IoT devices, leveraging malicious software (commonly known as malware) [3, 4]. Their motivations stem from several factors: first, the prevalence of legacy devices still connected to the Internet without adequate security updates; second, the tendency to prioritize other aspects over security during the development cycle; and third, the widespread use of weak login credentials. The sheer volume of deployed devices ranging from cameras to sensors, poses a formidable challenge when it comes to implementing cost-effective, real-time anomaly detection at such a massive scale.
@@
Securing IoT networks against intrusion is paramount, and network intrusion detection systems (NIDS) play a pivotal role in achieving this. Traditionally, NIDSs have relied heavily on signature-based methods [5, 6]- essentially matching attack patterns to known signatures. However, these approaches demand significant manual effort from experts to create and maintain those signatures. Moreover, they fall short when it comes to identifying novel, previously unseen attacks. With the introduction of machine learning (ML)-based NIDSs which leverage ML models to enhance existing defenses, impressive performance and accuracy has been achieved [7, 8]. Unlike their signature-centric counterparts, ML-based NIDSs do not require experts to painstakingly prepare signatures in advance. In fact, unsupervised ML-based NIDSs can even raise alerts for hitherto unknown attacks. Thanks to these advantages and the rapid evolution of ML techniques, ML-based NIDSs have made significant strides in recent years.
@@
Despite the promise of ML-based NIDSs, there is a noticeable hesitance among network security experts and practitioners when it comes to deploying them in real-world production environments [9]. The crux of the issue lies in the black-box nature of many of these proposed solutions. Essentially, their inner workings (how and why they make decisions) remain elusive, unlike simpler but less effective rule-based approaches that security professionals are accustomed to. Security analysts rely on ML-based NIDSs for critical decisions. However, without clear explanations for the decisions made by these underlying ML models, it becomes difficult for the analysts to trust them. Moreover, ML often operates as a black box within the NIDS. Consequently, hidden correlations stemming from unrelated artifacts can slip past undetected, leading to a flawed perception of the system’s capabilities. Explanation tools and techniques can unveil those hidden correlations, allowing practitioners to assess their impact on the NIDS’s overall effectiveness. As it stands, the current state-of-the-art NIDSs for IoT environments lack inherent explainability. Therefore, there is a pressing need to scrutinize them using these very explanation techniques.
@@
In this work, we analyze state-of-the-art ML-based IoT NIDS models using explainable AI (xAI) techniques (e.g., TRUSTEE [10], SHAP [11]). Except Kitsune, other NIDS source codes do not provide trained ML models or training/test data in a format which can be directly input to the xAI methods’ APIs. Therefore, we make changes in the NIDS source codes to satisfy the APIs’ requirements. Using both global explanations and local explanations for the models’ decisions, the most prominent features used by each ML-based IoT NIDS model considered are presented. We compare the explanations generated across xAI methods for a given NIDS model (using our devised qualitative approach) as well as the explanations generated across the NIDS models for a given xAI method (using standard evaluation metrics). Finally, we evaluate the vulnerability of each NIDS model to inductive bias (artifacts learnt from training data). To the best of our knowledge, our work is the first one to generate explanations for state-of-the-art ML-based IoT NIDSs other than Kitsune using xAI techniques.
@@
  • We analyze state-of-the-art ML-based IoT NIDS models using explainable AI techniques.
  • We present the most prominent features used by each IoT NIDS model considered to make decisions on input data samples using the explanations generated earlier.
@@
  • We evaluate the vulnerability of each IoT NIDS model to inductive bias.
The rest of the paper is organized as follows. In Section II, we review published research on the need for explanation of ML systems targeted at cybersecurity, evaluation of explanation methods on ML-based security systems and explainable ML-based attack detection. In Section III, we present a background on the types of explanation methods for ML models, prominent black-box xAI methods, and an overview of the state-of-the-art ML-based IoT NIDSs to which we apply explanation techniques in this work. In Section IV, we explain the issues with applying xAI methods to the ML-based IoT NIDSs directly using their open-source implementations, our approach to addressing those issues and details of our investigation beyond direct application of xAI methods which covers the relation between explanations generated by different xAI methods and the detection of inductive bias in the IoT NIDSs under consideration in this work. Finally, in Section V, we evaluate the explanations generated for each of the IoT NIDSs using xAI methods, compare the explanations across the NIDS models as well across xAI methods, and discuss insights gleaned from the main results obtained thus far.
@@ II Related Work
As a result, we can see the beginning of efforts in the research community to apply xAI techniques to ML-based security systems. A few works [14, 15, 16, 17] have employed xAI methods such as LIME [18] and SHAP [11] to assess the quality of local/global explanations generated for conventional machine learning classifiers (AdaBoost, k-nearest neighbour, Multi-layer Perceptron, Random Forest, XGBoost, Light Gradient-Boosting Machine, Gradient Boosting Classifier) applied to IoT botnet datasets (N-BaIoT [19], MedBIoT [20], and BoT-IoT [21]) and network intrusion datasets (NSL-KDD dataset [22], CICIDS-2017 [23]) in terms of criteria such as faithfulness, stability, complexity, and sensitivity. In [24], the authors have delved into six explanation methods, evaluating their effectiveness across four security systems described in existing literature. These systems leverage deep learning techniques to detect Android malware, malicious PDF files and security vulnerabilities. The assessment criteria encompass both general properties of deep learning and domain-specific aspects relevant to security.
@@
  • Existing works that evaluate xAI methods on ML-based IoT NIDSs do not focus on state-of-the-art IoT NIDSs and do not include the latest xAI tools for ML-based network security such as TRUSTEE ([14, 15, 16, 17]).
  • Existing explainable-by-design AI-based attack/intrusion detection systems are targeted at specific attacks (DNS-over-HTTPS attacks in [27]) and non-IoT networks (IT networks in [28]).
@@ III Background
In this paper, since we delve into explaining the decisions of state-of-the-art ML-based IoT NIDSs, it is essential to briefly explore the types of explanation methods present in academic literature, the IoT NIDSs that we consider for explanation in this work and the various aspects of our investigation with xAI techniques on those IoT NIDSs.
@@ III-B Prominent Black-box XAI Tools
ML-based IoT NIDS products are expected to be used by end users/customers who do not have access to the ML model and its parameters underlying the NIDS. This is to maintain confidentiality of the NIDS implementation by the cybersecurity services vendor. Therefore, we assume that the ML-based IoT NIDSs being explained in this work are available to us as a black-box system only and we use xAI techniques designed to explain such black-box ML systems. In what follows, we present a brief overview of each black-box xAI tool that we have tested with as part of this work.
@@
III-C State-of-the-art ML-based IoT NIDSs
We now present a brief overview of each state-of-the-art ML-based IoT NIDS that we investigate using xAI techniques in this work.
@@ III-C State-of-the-art ML-based IoT NIDSs
IV xAI Analysis of ML-based IoT NIDSs
In this work, we analyze four state-of-the-art ML-based IoT NIDSs (Kitsune, HorusEeye, ENIDrift, HELAD) using xAI techniques. However, xAI techniques such as TRUSTEE and kernel SHAP can not be applied directly to the IoT NIDSs under consideration. This is because xAI methods have certain requirements at the API level which are not fulfilled by the current implementations of the NIDSs. We had to modify the NIDS implementations to make them compatible with xAI methods. Still, there were a few xAI methods whose requirements were found to be incompatible with the NIDS implementations unless we made major changes to the source code. Additionally, we did not limit ourselves to a direct application of xAI methods to the NIDSs but we also investigated: (1) the relation between the explanations generated for a given NIDS by different xAI methods, and (2) the presence of inductive biases in the NIDS implementations.
IV-A xAI-compatible IoT NIDS Implementation
@@
While running LIME analysis of above NIDSs, we received an error that LIME does not support ML models which do not have $predict\_proba()$ function implemented. Since all the state-of-the-art ML-based IoT NIDSs considered in this paper do not implement $predict\_proba()$ function, we are forced to leave out LIME analysis of the NIDSs from this paper. Here, $predict\_proba()$ function returns the predicted probabilities of the input features belonging to each category. This method, instead of returning a discrete class, returns the probabilities associated with each class. This is useful when we not only want to know the predicted category of the input features, but also the model’s confidence in its prediction.
While running LEMNA analysis of above NIDSs, we found that the current implementation (last changed 5 years ago) only supports basic MLP, CNN and RNN models. Custom ML models such as the IoT NIDS models which are under consideration in this work are not supported. Further, LEMNA does not output feature-based model explanations for processing by cybersecurity experts such as those offered by TRUSTEE and SHAP.
@@ IV-B Investigation Beyond Direct xAI Analysis
Relation between TRUSTEE and SHAP Interpretations: To establish the relationship between TRUSTEE global explanations and the SHAP local explanations of the black-box model underlying a given IoT NIDS, we can follow two approaches:
@@
Inductive Bias: A number of ML models face the issue of underspecification. This term refers to the challenge of determining whether a model’s high performance is due to its genuine ability to capture essential patterns in the data or merely the result of certain inductive biases it has learned. These biases often appear as spurious correlations, a lack of ability to adapt to out-of-distribution (o.o.d.) data or shortcut learning strategies. The presence of such biases in trained models undermines their reliability and expected performance in real-world applications. Therefore, identifying these inductive biases is crucial for establishing trust in ML models. To detect the presence of an inductive bias in the ML-based IoT NIDSs under consideration in this work, such as vulnerability to o.o.d. samples, we use tampered malicious traffic traces [10] which alter the attack segments of the original traffic traces by distributing the malicious packets in such a way that the packet rate remains within a specified limit.
@@ V Experimental Results
In this section, we present explanations of the four state-of-the-art ML-based IoT NIDSs (Kitsune, HorusEeye, ENIDrift, HELAD) generated using xAI methods such as TRUSTEE and kernel SHAP. For each pair of NIDS and xAI method, we wrote Python scripts for generating the explanation using the xAI method’s API and the modified source code of the NIDS (to make it compatible with the xAI method’s API implementation). Using the generated explanations, we present the most prominent features used by each NIDS model to make decisions on input data samples. We also investigate the relation between the explanations generated for a given NIDS by TRUSTEE and kernel SHAP by using the systematic approach for comparing TRUSTEE and SHAP explanations at the DT subset-level as explained in sub-section IV-B. Subsequently, we investigate the presence of inductive biases (e.g., vulnerability to o.o.d. data samples) in the NIDS implementations using tampered malicious traffic traces as explained in sub-section IV-B. Finally, we compare the explanations generated across the four NIDSs by TRUSTEE using the fidelity metric and by kernel SHAP using general xAI evaluation metrics.
Datasets: We use the Kitsune Mirai dataset [32] and the CICIDS-2017 dataset [23] for training/testing the ML-based IoT NIDSs under consideration. The Mirai trace consists of $\sim$ 760k packets of synthetically generated attack in a network with nine IoT devices, first $\sim$120k packets consist of benign traffic, remaining $\sim$640k packets have anomalous traffic. The CICIDS-2017 dataset [23] consists of five days of packet capture (July 3 to July 7, 2017) from a small-scale enterprise-like network, with the first day consisting of benign background traffic and the other days consisting of attack traffic (brute force SSH, DoS, Heartbleed, web attack, infiltration, etc.). It consists of significant drift over the data collection period in terms of different attack tools used. Below, we present the results using Mirai dataset only. Since the results obtained with CICIDS-2017 dataset are similar to those obtained with the Mirai dataset, we include them in the Appendix Section A.

### v2

# Evaluating The Explainability of State-of-the-Art Machine Learning-based Online Network Intrusion Detection Systems

@@ Abstract
Network Intrusion Detection Systems (NIDSs) which use machine learning (ML) models achieve high detection performance and accuracy while avoiding dependence on fixed signatures extracted from attack artifacts. However, there is a noticeable hesitance among network security experts and practitioners when it comes to deploying ML-based NIDSs in real-world production environments due to their black-box nature, i.e., how and why the underlying models make their decisions. In this work, we analyze state-of-the-art ML-based online NIDS models using explainable AI (xAI) techniques (e.g., TRUSTEE, SHAP). Using the explanations generated for the models’ decisions, the most prominent features used by each NIDS model considered are presented. We compare the explanations generated across xAI methods for a given NIDS model as well as the explanations generated across the NIDS models for a given xAI method. Finally, we evaluate the vulnerability of each NIDS model to inductive bias (artifacts learnt from training data). The results show that: (1) some ML-based NIDS models can be better explained than other models, (2) xAI explanations are in conflict for most of the NIDS models considered in this work and (3) some NIDS models are more vulnerable to inductive bias than other models.
@@
Network Intrusion Detection Systems, NIDS, Machine Learning, explainable Artificial Intelligence, xAI
@@ I Introduction
Securing networks against intrusion is paramount, and network intrusion detection systems (NIDS) play a pivotal role in achieving this. Traditionally, NIDSs have relied heavily on signature-based methods [5, 6]- essentially matching attack patterns to known signatures. However, these approaches demand significant manual effort from experts to create and maintain those signatures. Moreover, they fall short when it comes to identifying novel, previously unseen attacks. With the introduction of machine learning (ML)-based NIDSs which leverage ML models to enhance existing defenses, impressive performance and accuracy has been achieved [7, 8]. Unlike their signature-centric counterparts, ML-based NIDSs do not require experts to painstakingly prepare signatures in advance. In fact, unsupervised ML-based NIDSs can even raise alerts for hitherto unknown attacks. Thanks to these advantages and the rapid evolution of ML techniques, ML-based NIDSs have made significant strides in recent years.
@@
Despite the promise of ML-based NIDSs, there is a noticeable hesitance among network security experts and practitioners when it comes to deploying them in real-world production environments [9]. The crux of the issue lies in the black-box nature of many of these proposed solutions. Essentially, their inner workings (how and why they make decisions) remain elusive, unlike simpler but less effective rule-based approaches that security professionals are accustomed to. Security analysts rely on ML-based NIDSs for critical decisions. However, without clear explanations for the decisions made by these underlying ML models, it becomes difficult for the analysts to trust them. Moreover, ML often operates as a black box within the NIDS. Consequently, hidden correlations stemming from unrelated artifacts can slip past undetected, leading to a flawed perception of the system’s capabilities. Explanation tools and techniques can unveil those hidden correlations, allowing practitioners to assess their impact on the NIDS’s overall effectiveness. As it stands, the current state-of-the-art online NIDSs lack inherent explainability. Therefore, there is a pressing need to scrutinize them using these very explanation techniques.
@@
In this work, we analyze state-of-the-art ML-based online NIDS models using explainable AI (xAI) techniques (e.g., TRUSTEE [10], SHAP [11]). Except Kitsune, other NIDS source codes do not provide trained ML models or training/test data in a format which can be directly input to the xAI methods’ APIs. Therefore, we make changes in the NIDS source codes to satisfy the APIs’ requirements. Using both global explanations and local explanations for the models’ decisions, the most prominent features used by each ML-based NIDS model considered are presented. We compare the explanations generated across xAI methods for a given NIDS model (using our devised qualitative approach) as well as the explanations generated across the NIDS models for a given xAI method (using standard evaluation metrics). Finally, we evaluate the vulnerability of each NIDS model to inductive bias (artifacts learnt from training data). To the best of our knowledge, our work is the first one to generate explanations for state-of-the-art ML-based online NIDSs other than Kitsune using xAI techniques.
@@
  • We analyze state-of-the-art ML-based online NIDS models using explainable AI techniques.
  • We present the most prominent features used by each NIDS model considered to make decisions on input data samples using the explanations generated earlier.
@@
  • We evaluate the vulnerability of each NIDS model to inductive bias.
The rest of the paper is organized as follows. In Section II, we review published research on the need for explanation of ML systems targeted at cybersecurity, evaluation of explanation methods on ML-based security systems and explainable ML-based attack detection. In Section III, we present a background on the types of explanation methods for ML models, prominent black-box xAI methods, and an overview of the state-of-the-art ML-based online NIDSs to which we apply explanation techniques in this work. In Section IV, we explain the issues with applying xAI methods to the ML-based NIDSs directly using their open-source implementations, our approach to addressing those issues and details of our investigation beyond direct application of xAI methods which covers the relation between explanations generated by different xAI methods and the detection of inductive bias in the online NIDSs under consideration in this work. Finally, in Section V, we evaluate the explanations generated for each of the NIDSs using xAI methods, compare the explanations across the NIDS models as well across xAI methods, and discuss insights gleaned from the main results obtained thus far.
@@ II Related Work
As a result, we can see the beginning of efforts in the research community to apply xAI techniques to ML-based security systems. A few works [14, 15, 16, 17] have employed xAI methods such as LIME [18] and SHAP [11] to assess the quality of local/global explanations generated for conventional machine learning classifiers (AdaBoost, k-nearest neighbour, Multi-layer Perceptron, Random Forest, XGBoost, Light Gradient-Boosting Machine, Gradient Boosting Classifier) applied to network intrusion datasets (NSL-KDD dataset [22], CICIDS-2017 [23]) in terms of criteria such as faithfulness, stability, complexity, and sensitivity. In [24], the authors have delved into six explanation methods, evaluating their effectiveness across four security systems described in existing literature. These systems leverage deep learning techniques to detect Android malware, malicious PDF files and security vulnerabilities. The assessment criteria encompass both general properties of deep learning and domain-specific aspects relevant to security.
@@
  • Existing works that evaluate xAI methods on ML-based NIDSs do not focus on state-of-the-art online NIDSs and do not include the latest xAI tools for ML-based network security such as TRUSTEE ([14, 15, 16, 17]).
  • Existing explainable-by-design AI-based attack/intrusion detection systems are targeted at specific attacks (DNS-over-HTTPS attacks in [27]).
@@ III Background
In this paper, since we delve into explaining the decisions of state-of-the-art ML-based online NIDSs, it is essential to briefly explore the types of explanation methods present in academic literature, the NIDSs that we consider for explanation in this work and the various aspects of our investigation with xAI techniques on those NIDSs.
@@ III-B Prominent Black-box XAI Tools
ML-based NIDS products are expected to be used by end users/customers who do not have access to the ML model and its parameters underlying the NIDS. This is to maintain confidentiality of the NIDS implementation by the cybersecurity services vendor. Therefore, we assume that the ML-based NIDSs being explained in this work are available to us as a black-box system only and we use xAI techniques designed to explain such black-box ML systems. In what follows, we present a brief overview of each black-box xAI tool that we have tested with as part of this work.
@@
III-C State-of-the-art ML-based Online NIDSs
We now present a brief overview of each state-of-the-art ML-based online NIDS that we investigate using xAI techniques in this work.
@@ III-C State-of-the-art ML-based Online NIDSs
IV xAI Analysis of ML-based Online NIDSs
In this work, we analyze four state-of-the-art ML-based online NIDSs (Kitsune, HorusEeye, ENIDrift, HELAD) using xAI techniques. However, xAI techniques such as TRUSTEE and kernel SHAP can not be applied directly to the NIDSs under consideration. This is because xAI methods have certain requirements at the API level which are not fulfilled by the current implementations of the NIDSs. We had to modify the NIDS implementations to make them compatible with xAI methods. Still, there were a few xAI methods whose requirements were found to be incompatible with the NIDS implementations unless we made major changes to the source code. Additionally, we did not limit ourselves to a direct application of xAI methods to the NIDSs but we also investigated: (1) the relation between the explanations generated for a given NIDS by different xAI methods, and (2) the presence of inductive biases in the NIDS implementations.
IV-A xAI-compatible Online NIDS Implementation
@@
While running LIME analysis of above NIDSs, we received an error that LIME does not support ML models which do not have $predict\_proba()$ function implemented. Since all the state-of-the-art ML-based online NIDSs considered in this paper do not implement $predict\_proba()$ function, we are forced to leave out LIME analysis of the NIDSs from this paper. Here, $predict\_proba()$ function returns the predicted probabilities of the input features belonging to each category. This method, instead of returning a discrete class, returns the probabilities associated with each class. This is useful when we not only want to know the predicted category of the input features, but also the model’s confidence in its prediction.
While running LEMNA analysis of above NIDSs, we found that the current implementation (last changed 5 years ago) only supports basic MLP, CNN and RNN models. Custom ML models such as the NIDS models which are under consideration in this work are not supported. Further, LEMNA does not output feature-based model explanations for processing by cybersecurity experts such as those offered by TRUSTEE and SHAP.
@@ IV-B Investigation Beyond Direct xAI Analysis
Relation between TRUSTEE and SHAP Interpretations: To establish the relationship between TRUSTEE global explanations and the SHAP local explanations of the black-box model underlying a given NIDS, we can follow two approaches:
@@
Inductive Bias: A number of ML models face the issue of underspecification. This term refers to the challenge of determining whether a model’s high performance is due to its genuine ability to capture essential patterns in the data or merely the result of certain inductive biases it has learned. These biases often appear as spurious correlations, a lack of ability to adapt to out-of-distribution (o.o.d.) data or shortcut learning strategies. The presence of such biases in trained models undermines their reliability and expected performance in real-world applications. Therefore, identifying these inductive biases is crucial for establishing trust in ML models. To detect the presence of an inductive bias in the ML-based NIDSs under consideration in this work, such as vulnerability to o.o.d. samples, we use tampered malicious traffic traces [10] which alter the attack segments of the original traffic traces by distributing the malicious packets in such a way that the packet rate remains within a specified limit.
@@ V Experimental Results
In this section, we present explanations of the four state-of-the-art ML-based online NIDSs (Kitsune, HorusEeye, ENIDrift, HELAD) generated using xAI methods such as TRUSTEE and kernel SHAP. For each pair of NIDS and xAI method, we wrote Python scripts for generating the explanation using the xAI method’s API and the modified source code of the NIDS (to make it compatible with the xAI method’s API implementation). Using the generated explanations, we present the most prominent features used by each NIDS model to make decisions on input data samples. We also investigate the relation between the explanations generated for a given NIDS by TRUSTEE and kernel SHAP by using the systematic approach for comparing TRUSTEE and SHAP explanations at the DT subset-level as explained in sub-section IV-B. Subsequently, we investigate the presence of inductive biases (e.g., vulnerability to o.o.d. data samples) in the NIDS implementations using tampered malicious traffic traces as explained in sub-section IV-B. Finally, we compare the explanations generated across the four NIDSs by TRUSTEE using the fidelity metric and by kernel SHAP using general xAI evaluation metrics.
Datasets: We use the Kitsune Mirai dataset [32] and the CICIDS-2017 dataset [23] for training/testing the ML-based NIDSs under consideration. The Mirai trace consists of $\sim$ 760k packets of synthetically generated attack in a network with nine IoT devices, first $\sim$120k packets consist of benign traffic, remaining $\sim$640k packets have anomalous traffic. The CICIDS-2017 dataset [23] consists of five days of packet capture (July 3 to July 7, 2017) from a small-scale enterprise-like network, with the first day consisting of benign background traffic and the other days consisting of attack traffic (brute force SSH, DoS, Heartbleed, web attack, infiltration, etc.). It consists of significant drift over the data collection period in terms of different attack tools used. Below, we present the results using Mirai dataset only. Since the results obtained with CICIDS-2017 dataset are similar to those obtained with the Mirai dataset, we include them in the Appendix Section A.